book 31 i got this equation wrong. Rewrite the equation –8x – 5y = –45 as a function of x.
Answer by mszlmb(115)

"Rewrite the equation –8x – 5y = –45 as a function of x."

-8x-5y=-45+5y on both sides
-8x=5y-45divide by -8
x=-5/8y+45/8OPTIONAL: Clean it up:
x=5.625-.625yOR x=5/8-5 5/8y